OpenStudy (lgbasallote):

how can this be factored? \[\Large \text D^3 - 11\text D - 20\] is it different from algebra?

OpenStudy (lgbasallote):

is the factoring in diff eq different from algebra?

OpenStudy (anonymous):

no

OpenStudy (lgbasallote):

so how do you factor diff eqs?

OpenStudy (anonymous):

Rational root thereom :)

OpenStudy (lgbasallote):

but isnt that how you factor algebra as well? what's the difference?

OpenStudy (anonymous):

I am thinking if it is factorizable..

OpenStudy (lgbasallote):

my first question is if factoring diff eqs is different from algebra

OpenStudy (anonymous):

It is somewhat different...

OpenStudy (callisto):

Let f(D) = D^3 -11D - 20 f(4) = 4^3-11(4) -20 = 0 Therefore, (x-4) is a factor of f(x) Use synthetic division.  f(x) = (x-4)(x^2+4x+5)